# Improving Efficiency of Diffusion Models via Multi-Stage Framework and Tailored Multi-Decoder Architectures

### v1

@@ 1 Introduction
Building on these insights, we propose a multi-stage framework with two key components: (1) a multi-decoder U-net architecture, and (2) a new partitioning algorithm to cluster timesteps (noise levels) into distinct stages. In terms of our new architecture, we design a multi-decoder U-Net that incorporates one universal encoder shared across all intervals and individual decoders tailored to each time stage; see Figure 1 (c) for an illustration. This approach combines the advantages of both universal and stage-specific architectures, which is much more efficient that using a single architecture for the entire training process [2, 1, 18] (Figure 1 (a)). Moreover, compared to previous approaches that completely separate architectures for each sub-interval [19, 20, 21, 22] (Figure 1 (b)), our approach can effectively mitigate interference between stages arising from disparate gradient effects, leading to improved efficiency. On the other hand, when it comes to partitioning the training stages of our network, we designed an algorithm aimed at grouping the timesteps. This is achieved by minimizing the functional distance within each cluster in the training objective and making use of the optimal denoiser formulation [18]. By integrating these two key components, our framework enables efficient allocation of computational resources (e.g., U-net parameters) and stage-tailored parameterization. In Section 5, we showcase via extensive numerical experiments that our framework effectively improves both training and sampling efficiency. These experiments are performed on diverse benchmark datasets, demonstrating significant acceleration by using our framework when compared to three state-of-the-art (SOTA) diffusion model architectures.
@@
  • Identifying two key sources of inefficiency. We identified two key sources that cause inefficiencies in training diffusion models across various time step: (1) a significant variation in the requirement of model capacity, and (2) the gradient dissimilarity. As such, using a unified network cannot meet with the changing requirement at different time steps.
  • A new multi-stage framework. We introduced a new multi-stage architecture, illustrated in Figure 1 (c). We tackle the challenges by segmenting the time interval into multiple stages, where we employ customized multi-decoder U-net architectures that blends time-dependent models with a universally shared encoder.
  • Improved training and sampling efficiency. With comparable computational resources for unconditional image generation, we demonstrate that our multi-stage approach improves the Fréchet Inception Distance (FID) score for all SOTA methods. For example, on CIFAR-10 dataset [23], our method improves the FID for DPM-Solver [24] from 2.84 to 2.37, and it improves the FID for EDM [18] from 2.05 (our training result) to 1.96. Moreover, on the CelebA dataset [25], our approach significantly reduces the training computation of both EDM to 82% and the Latent Diffusion Model (LDM) [8] to 30% for obtaining a similar quality of generation.
@@
In Section 2, we provide an overview of relevant literature to distinguish us from previous works. In Section 3, we describe our proposed multistage framework for diffusion modeling, delineating the two core components. In Section 4, we present our observations and analysis that motivated the proposed multistage framework, justifying its development. Finally, in Section 5, we provide the results from our numerical experiments that validate the effectiveness of the proposed multistage approach.
@@ 2 Preliminaries & Related Work
In this section, we start by reviewing the basic fundamentals of diffusion models [1, 2, 18]. Subsequently, we delve into prior endeavors aimed at improving the training and efficiency of diffusion models through the partitioning of the timestep interval. Lastly, we revisit prior studies that significantly decrease the number of required sampling iterations.
2.1 Background on Diffusion Models
Let $\bm{x}_{0}\in\mathbb{R}^{n}$ denote a sample from the data distribution $p_{\text{data}}(\bm{x})$. Diffusion models operate within forward and reverse processes. The forward process progressively perturbs data $\bm{x}_{0}$ to a noisy version $\bm{x}_{t\in[0,1]}$ via corrupting with the Gaussian kernel. This process can be formulated as a stochastic differential equation (SDE)  [2] of the form $\text{d}\bm{x}=\bm{x}_{t}f(t)\text{d}t+g(t)\text{d}\bm{w}_{t}$, where $f(t)$ and $g(t)$ are the drift and diffusion coefficients, respectively, that correspond to a pre-defined noise schedule. $\bm{w}_{t}\in\mathbb{R}^{n}$ is the standard Wiener process. Under the forward SDE, the perturbation kernel is given by the conditional distribution defined as
$p_{t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I})\>,$ (1)
where
@@ 2.1 Background on Diffusion Models
The parameters defined in eq. 2 are designed such that: (i) the data distribution is approximately estimated when $t=0$, and (ii) a nearly standard Gaussian distribution is obtained when $t=1$. The objective of diffusion models is to learn the corresponding reverse SDE, defined as
$\text{d}\bm{x}=\left[f(t)\bm{x}_{t}-g^{2}(t)\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})\right]\text{d}t+g(t)\text{d}\bar{\bm{w}},$ (3)
where $\bar{\bm{w}}\in\mathbb{R}^{n}$ is the standard Wiener process running backward in time, and $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ is the (Stein) score function. In practice, the score function is approximated using a neural network $\bm{\epsilon}_{\bm{\theta}}:\mathbb{R}^{n}\times[0,1]\rightarrow\mathbb{R}^{n}$ parameterized by $\bm{\theta}$, which can be trained by the denoising score matching technique [26] as
@@ 2.2 Timestep Clustering
Diffusion models have demonstrated exceptional performance but face efficiency challenges in training and sampling. In response, several studies proposed to cluster the timestep range $t\in[0,1]$ into multiple intervals (e.g., $[0,t_{1}),[t_{1},t_{2}),\ldots,[t_{n},1]$). Notably, Choi et al. [19] reconfigured the loss weights for different intervals to enhance performance. Balaji et al. [27] introduced “expert denoisers”, which proposed using distinct architectures for different time intervals in text-to-image diffusion models. Go et al. [22] further improved the efficiency of these expert denoisers through parameter-efficient fine-tuning and data-free knowledge transfer. Lee et al. [21] designed separate architectures for each interval based on frequency characteristics. Moreover, Go et al. [20] treated different intervals as distinct tasks and employed multi-task learning strategies for diffusion model training, along with various timestep clustering methods.
@@
Our approach distinguishes itself from the aforementioned methods in two key aspects. The first key component is our tailored U-net architecture using a unified encoder coupled with different decoders for different intervals, while previous models have either adopted a unified architecture, as seen in [19, 20], or employed separate architectures (referred to as expert denoisers) for each interval [27, 22, 21]. In comparison, our multistage architecture surpasses these methodologies, as demonstrated in Section 5.3. Second, we developed a new timestep clustering method leveraging a general optimal denoiser (Proposition 1) that showcases superior performance (see Section 5.4). In contrast, prior works rely on (i) a simple timestep-based clustering cost function [27, 20, 22, 21], (ii) Signal-to-Noise Ratio (SNR) based clustering [20], or (iii) gradients-based partitioning that uses task affinity scores [20].
2.3 Reducing the Sampling Iterations
@@
Efforts to improve sampling efficiency of diffusion models have led to many recent advancements in SDE and Ordinary Differential Equation (ODE) samplers [2]. For instance, the Denoising Diffusion Implicit Model (DDIM) [28] formulates the forward diffusion as a non-Markovian process with a deterministic generative path, significantly reducing the number of function evaluations (NFE) required for sampling (from thousands to hundreds) without sacrificing generation quality. Generalized DDIM (gDDIM) [29] further optimized DDIM by modifying the parameterization of the scoring network, reducing NFE to nearly 27. Furthermore, the works in [24] and [30], termed the Diffusion Probabilistic Model solver (DPM-solver) and the Diffusion Exponential Integrator Sampler (DEIS), respectively, introduced fast higher-order solvers, employing exponential integrators that require even less than 20 NFE for comparable generation quality. Moreover, the consistency model [31] introduced a novel training loss and parameterization, achieving high-quality generation with merely 1-2 NFE.
@@ 3 Proposed Multistage Framework
In this section, we introduce our new multistage framework (as illustrated in Figure 1 (c)). Specifically, we first introduce the multi-stage U-Net architecture design in Section 3.1, following a new clustering method for choosing the optimal interval check points to partition the entire timestep $[0,1]$ into intervals in Section 3.2. We leave the discussion on the motivation behind our method to Section 4.
@@ 3.1 Proposed Multi-stage U-Net Architectures
Most existing diffusion models either employ a unified architecture across all intervals [1, 2, 18] to share features for all timesteps, or they use completely separate architectures for different timestep intervals [21, 22, 27] to take advantage of benign properties within different intervals. The details of the advantages and disadvantages of each architecture are discussed in section 4.
To harness the advantages of both unified and separate architectures employed in prior studies, we introduce a multistage U-Net architecture, as illustrated in Figure 1(c). Specifically, we partition the entire timestep $[0,1]$ into several intervals, e.g., three intervals $[0,t_{1}),[t_{1},t_{2}),[t_{2},1]$ in Figure 1. For the architecture, we introduce:
  • One shared encoder across all time intervals. For each timestep interval, we implement a shared encoder architecture (plotted in blue in Figure 1 (c)), which is similar to the architecture employed in the original U-Net framework [32]. Unlike separate architecture, the shared encoder provide shared information across all timesteps, preventing models from overfitting (see Section 5.3 for a discussion).
  • Separate decoders for different time intervals. Motivated by the multi-head structure introduced in the Mask Region-based Convolutional Neural Networks (Mask-RCNN) method [33], we propose to use multiple distinct decoders (plotted in colors for different intervals in Figure 1(c)), where each decoder is tailored to a specific timestep interval. The architecture of each decoder closely resembles the one utilized in [2], with deliberate adjustments made to the embedding dimensions to optimize performance.
As we observe, the primary difference in the architecture resides within the decoder structure. Intuitively, we use a decoder with fewer number of parameters for intervals closer to the noise, because the learning task is easier. We use a decoder with larger number of parameters for intervals closer to the image. In Section 4, we conduct a comprehensive analysis to elucidate the rationale behind the adoption of this multistage architecture, and the necessity for varying parameters across different intervals.
@@ 3.2 Optimal Denoiser-based Timestep Clustering
Next, we discuss how we principally choose the interval partition time points in practice. For simplicity, we focus on the case where we partition the time $[0,1]$ into three intervals $[0,t_{1}),[t_{1},t_{2}),[t_{2},1]$, and we develop a timestep clustering method to choose the optimal $t_{1}$ and $t_{2}$ that we introduce in the following. Of course, our method can be generalized to multi-stage networks with arbitrary interval numbers. However, in practice, we find that the choice of three intervals strikes a good balance between effectiveness and complexity; see our ablation study in the appendices.
To partition the time interval, we employ the optimal denoiser established in Proposition 1, where we can show the following result.
@@
1:Input: Total samples $K$, optimal denoiser function $\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x},t)$, thresholds $\alpha$, $\eta$, dataset $p_{\text{data}}$
@@
3:$\mathcal{S}_{0}=\mathcal{S}_{1}=\emptyset$
4:for $k\in\{1,\dots,K\}$ do
5:     $\bm{y}_{k}\sim p_{\text{data}},\bm{\epsilon}_{k}\sim\mathcal{N}(0,\mathbf{I}),t_{k}\sim[0,1]$
6:     $\mathcal{S}^{k}_{0}\leftarrow\mathcal{D}(\bm{\epsilon}^{*}_{t_{k}},\bm{\epsilon}^{*}_{0},\bm{y}_{k},\bm{\epsilon}_{k}),\mathcal{S}^{k}_{1}\leftarrow\mathcal{D}(\bm{\epsilon}^{*}_{t_{k}},\bm{\epsilon}^{*}_{1},\bm{y}_{k},\bm{\epsilon}_{k})$
7:     $\mathcal{S}_{0}\leftarrow\mathcal{S}_{0}\cup\{\left(t_{k},\mathcal{S}^{k}_{0}\right)\}$
8:     $\mathcal{S}_{1}\leftarrow\mathcal{S}_{1}\cup\{\left(t_{k},\mathcal{S}^{k}_{1}\right)\}$
9:end for
10:$t_{1}=\operatorname*{\text{arg~{}max}}\limits_{\tau}\left\{\tau~{}\Big{|}~{}\dfrac{\sum_{(t_{k},\mathcal{S}^{k}_{0})\in\mathcal{S}_{0}}\left[\mathcal{S}^{k}_{0}\cdot\mathds{1}(t_{k}\leq\tau)\right]}{\sum_{(t_{k},\mathcal{S}^{k}_{0})\in\mathcal{S}_{0}}\left[\mathds{1}(t_{k}\leq\tau)\right]}\geq\alpha\right\}$
11:$t_{2}=\operatorname*{\text{arg~{}min}}\limits_{\tau}\left\{\tau~{}\Big{|}~{}\dfrac{\sum_{(t_{k},\mathcal{S}^{k}_{1})\in\mathcal{S}_{0}}\left[\mathcal{S}^{k}_{1}\cdot\mathds{1}(t_{k}\geq\tau)\right]}{\sum_{(t_{k},\mathcal{S}^{k}_{1})\in\mathcal{S}_{0}}\left[\mathds{1}(t_{k}\geq\tau)\right]}\geq\alpha\right\}$
@@
[…]
$\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)=\frac{1}{s_{t}\sigma_{t}}\left[\bm{x}-s_{t}\frac{\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})\bm{y}_{i}}{\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})}\right]\>.$
The proof is provided in appendices, and the result can be generalized from recent work of Karras et al. [18], extending from a specific kernel $p_{t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\bm{x}_{0},\sigma_{t}^{2}\textbf{I})$ to encompassing a broader scope of noise perturbation kernels, given by $p_{t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I})$. For brevity, we simplify the notation of the optimal denoiser $\bm{\epsilon}^{*}_{\theta}(\bm{x},t)$ in Proposition 1 as $\bm{\epsilon}^{*}_{t}(\bm{x})$.

### v2

@@ 1 Introduction
Building on these insights, we propose a multi-stage framework with two key components: (1) a multi-decoder U-net architecture, and (2) a new partitioning algorithm to cluster timesteps (noise levels) into distinct stages. In terms of our new architecture, we design a multi-decoder U-Net that incorporates one universal encoder shared across all intervals and individual decoders tailored to each time stage; see fig. 1 (c) for an illustration. This approach combines the advantages of both universal and stage-specific architectures, which is much more efficient that using a single architecture for the entire training process [2, 1, 18] (fig. 1 (a)). Moreover, compared to previous approaches that completely separate architectures for each sub-interval [19, 20, 21, 22] (fig. 1 (b)), our approach can effectively mitigate interference between stages arising from disparate gradient effects, leading to improved efficiency. On the other hand, when it comes to partitioning the training stages of our network, we designed an algorithm aimed at grouping the timesteps. This is achieved by minimizing the functional distance within each cluster in the training objective and making use of the optimal denoiser formulation [18]. By integrating these two key components, our framework enables efficient allocation of computational resources (e.g., U-net parameters) and stage-tailored parameterization. In section 5, we showcase via extensive numerical experiments that our framework effectively improves both training and sampling efficiency. These experiments are performed on diverse benchmark datasets, demonstrating significant acceleration by using our framework when compared to three state-of-the-art (SOTA) diffusion model architectures.
@@
  • Identifying two key sources of inefficiency. We identified two key sources that cause inefficiencies in training diffusion models across various time step: (i) a significant variation in the requirement of model capacity, and (ii) the gradient dissimilarity. As such, using a unified network cannot meet with the changing requirement at different time steps.
  • A new multi-stage framework. We introduced a new multi-stage architecture, illustrated in fig. 1 (c). We tackle these two sources of inefficiency by segmenting the time interval into multiple stages, where we employ customized multi-decoder U-net architectures that blends time-dependent models with a universally shared encoder.
  • Improved training and sampling efficiency. With comparable computational resources for unconditional image generation, we demonstrate that our multi-stage approach improves the Fréchet Inception Distance (FID) score for all SOTA methods. For example, on CIFAR-10 dataset [23], our method improves the FID for DPM-Solver [24] from 2.84 to 2.37, and it improves the FID for EDM [18] from 2.05 (our training result) to 1.96. Moreover, on the CelebA dataset [25], while maintaining a similar generation quality, our approach significantly reduces the required training FLOPS of EDM by 82% and the Latent Diffusion Model (LDM) [8] by 30%.
@@
In section 2, we provide preliminaries and an overview of related literatures. In section 3, we present our observations and analysis that motivated the proposed multistage framework, justifying its development. In section 4, we describe our proposed multistage framework for diffusion models, outlining the two core components. Finally, in section 5, we provide the results from our numerical experiments that validate the effectiveness of the proposed multistage approach.
@@ 2 Preliminaries & Related Work
In this section, we start by reviewing the basic fundamentals of diffusion models [1, 2, 18]. Subsequently, we delve into prior approaches aimed at improving the training and efficiency of diffusion models through the partitioning of the timestep interval. Lastly, we survey prior studies that significantly decrease the number of required sampling iterations.
Background on diffusion models. Let $\bm{x}_{0}\in\mathbb{R}^{n}$ denote a sample from the data distribution $p_{\text{data}}(\bm{x})$. Diffusion models operate within forward and reverse processes. The forward process progressively perturbs data $\bm{x}_{0}$ to a noisy version $\bm{x}_{t\in[0,1]}$ via corrupting with the Gaussian kernel. This process can be formulated as a stochastic differential equation (SDE)  [2] of the form $\text{d}\bm{x}_{t}=\bm{x}_{t}f(t)\text{d}t+g(t)\text{d}\bm{w}_{t}$, where $f(t)$ and $g(t)$ are the drift and diffusion coefficients, respectively, that correspond to a pre-defined noise schedule. $\bm{w}_{t}\in\mathbb{R}^{n}$ is the standard Wiener process. Under the forward SDE, the perturbation kernel is given by the conditional distribution defined as $p_{t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I}),$ where
@@
The parameters $s_{t}$ and $\sigma_{t}$ are designed such that: (i) the data distribution is approximately estimated when $t=0$, and (ii) a nearly standard Gaussian distribution is obtained when $t=1$. The objective of diffusion models is to learn the corresponding reverse SDE, defined as $\text{d}\bm{x}=\left[f(t)\bm{x}_{t}-g^{2}(t)\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})\right]\text{d}t+g(t)\text{d}\bar{\bm{w}},$ where $\bar{\bm{w}}\in\mathbb{R}^{n}$ is the standard Wiener process running backward in time, and $\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})$ is the (Stein) score function. In practice, the score function is approximated using a neural network $\bm{\epsilon}_{\bm{\theta}}:\mathbb{R}^{n}\times[0,1]\rightarrow\mathbb{R}^{n}$ parameterized by $\bm{\theta}$, which can be trained by the denoising score matching technique [26] as
@@
Timestep clustering methods. Diffusion models have demonstrated exceptional performance but face efficiency challenges in training and sampling. In response, several studies proposed to cluster the timestep range $t\in[0,1]$ into multiple intervals (e.g., $[0,t_{1}),[t_{1},t_{2}),\ldots,[t_{n},1]$). Notably, Choi et al. [19] reconfigured the loss weights for different intervals to enhance performance. Deja et al. [27] divide the entire process into a denoiesr and a generator based on their functionalities. Balaji et al. [28] introduced “expert denoisers”, which proposed using distinct architectures for different time intervals in text-to-image diffusion models. Go et al. [22] further improved the efficiency of these expert denoisers through parameter-efficient fine-tuning and data-free knowledge transfer. Lee et al. [21] designed separate architectures for each interval based on frequency characteristics. Moreover, Go et al. [20] treated different intervals as distinct tasks and employed multi-task learning strategies for diffusion model training, along with various timestep clustering methods.
@@
Our approach distinguishes itself from the aforementioned methods in two key aspects. The first key component is our tailored U-net architecture using a unified encoder coupled with different decoders for different intervals. Previous models have either adopted a unified architecture, as seen in [19, 20], or employed separate architectures (referred to as expert denoisers) for each interval [28, 22, 21]. In comparison, our multistage architecture surpasses these methodologies, as demonstrated in section 5.3. Second, we developed a new timestep clustering method leveraging a general optimal denoiser (Prop. 1) that showcases superior performance (see section 5.4). In contrast, prior works rely on (i) a simple timestep-based clustering cost function [28, 20, 22, 21], (ii) Signal-to-Noise Ratio (SNR) based clustering [20], or (iii) gradients-based partitioning that uses task affinity scores [20].
@@
Reducing the sampling iterations methods. Efforts to improve sampling efficiency of diffusion models have led to many recent advancements in SDE and Ordinary Differential Equation (ODE) samplers [2]. For instance, the Denoising Diffusion Implicit Model (DDIM) [29] formulates the forward diffusion as a non-Markovian process with a deterministic generative path, significantly reducing the number of function evaluations (NFE) required for sampling (from thousands to hundreds). Generalized DDIM (gDDIM) [30] further optimized DDIM by modifying the parameterization of the scoring network. Furthermore, the works in [24] and [31], termed the Diffusion Probabilistic Model solver (DPM-solver) and the Diffusion Exponential Integrator Sampler (DEIS), respectively, introduced fast higher-order solvers, employing exponential integrators that require 10s NFE for comparable generation quality. Moreover, the consistency model [32] introduced a novel training loss and parameterization, achieving high-quality generation with merely 1-2 NFE.
@@
3 Identification of Key Sources of Inefficiency
Conventional diffusion model architectures, as exemplified by [2, 1, 18], treat the training of the diffusion model as a unified process across all timesteps. Recent researches, such as [19, 20, 21, 22], have highlighted the benefits of recognizing distinctions between different timesteps and the potential efficiency gains from treating them as separate tasks during the training process. However, our experimental results demonstrate that both unified and separate architectures suffer inefficiency for training diffusion models, where the inefficiency comes from (i) overparameterization, (ii) gradient dissimilarity, and (iii) overfitting.
3.1 Empirical Observations on the Key Sources of Inefficiency
(a) Analysis on interval $[0,t_{1})$
(b) Analysis on interval $[t_{2},1]$
Experiment setup.
In our experiments, we consider three-stage training and divide the time steps into three intervals: $[0,t_{1}),[t_{1},t_{2}),[t_{2},1]$ [^3] Details for interval clustering can be found in section 5.. Let $(\bm{\epsilon}_{\bm{\theta}})^{[a,b]}_{i},0\leq a<b\leq 1$ denote a U-Net architecture with parameter $\bm{\theta}$ trained with $i$ iterations and fed with data pairs $(\bm{x}_{t},t)$, where $t\in[a,b]$. We then train models using two different strategies: a unified architecture with 108M network parameters for all intervals, i.e., $(\bm{\epsilon}_{\bm{\theta}})^{[0,1]}_{i}$, and separate architectures with varying network parameters (e.g., 47M, 108M, 169M) for each interval; e.g., $(\bm{\epsilon}_{\bm{\theta}})^{[0,t_{1})}_{i}$ for the interval $[0,t_{1})$, etc. It is worth noting that, apart from the differences in network parameters, we utilize the same network architecture (e.g., U-Net) for both the unified and the separate approaches. We assessed the training progress of each model by evaluating image generation quality at different training iterations. Notably, because some of the models are only trained on one interval, we need to provide a ground truth score for the other intervals. In fig. 2, the sampling process is shown top and the experimental results are shown bottom.
Inefficiency in unified architectures.
From fig. 2, we observe the following:
  • Overparameterization and underfitting emerge simultaneously for unified architectures. In fig. 2(a), we observe that increasing the number of parameters in Interval 0 can improve the image generation quality (as indicated by a lower FID score). In contrast, fig. 2(b) reveals that increasing the number of parameters in Interval 2 has minimal impact on the quality of image generation. This implies that using a unified architecture will result in underfitting in Interval 0 and overparameterization in Interval 2. The current unified architecture’s parameter redundancy leaves significant room for improving its efficiency. To optimize the computational usage, we should allocate more parameters to Interval 0 while allocating fewer parameters to Interval 2.
  • Gradient dissimilarity hinders training for unified architecture. Quantitative results from [20] demonstrate dissimilarity in gradients caused by differences within each interval. This can also be observed from our results based upon both fig. 2(a) and fig. 2(b). For the unified and separate architectures using the same number of parameters (108M), separate architecture achieves a significantly lower FID with the same training iterations, implying that dissimilar gradients among intervals may hinder training when using a unified architecture. Here, the only difference between training separate and unified architectures is that the batch gradient for unified architecture is calculated based on all timesteps while the batch gradien2 for separate architecture is calculated only from a specific interval.
Inefficiency in existing separate architectures.
Although separated architecture [19, 21, 22] better allocates computational resources for each interval, it suffers from overfitting. This can be illustrated based upon training separate architectures (169M) and (108M) in Interval 0 shown in fig. 2(a), where increasing the number of parameters will lead to overfitting. This also happens in Interval 2, when we compare all separate architectures in fig. 2(b). In comparison, the unified networks with 108M parameters are less prone to overfitting for both Interval 0 and Interval 2. This suggests that we can reduce overfitting by training shared weights across different intervals together.
3.2 Tackling the Inefficiency via Multistage U-Net Architectures
@@ 4 Proposed Multistage Framework
In this section, we introduce our new multistage framework (as illustrated in fig. 1 (c)). Specifically, we first introduce the multi-stage U-Net architecture design in section 4.1, following a new clustering method for choosing the optimal interval to partition the entire timestep $[0,1]$ into intervals in section 4.2. , and discuss the rationale of the proposed architecture in section 4.3.
@@ 4.1 Proposed Multi-stage U-Net Architectures
As discussed in section 3, most existing diffusion models either employ a unified architecture across all intervals [1, 2, 18] to share features for all timesteps, or use completely separate architectures for different timestep intervals [21, 22, 28] where the goal is to take advantage of the benign properties within different intervals.
To harness the advantages of both unified and separate architectures employed in prior studies, we introduce a multistage U-Net architecture, as illustrated in Figure 1(c). Specifically, we partition the entire timestep $[0,1]$ into several intervals, e.g., three intervals $[0,t_{1}),[t_{1},t_{2}),[t_{2},1]$ in fig. 1. For the architecture, we introduce:
  • One shared encoder across all time intervals. For each timestep interval, we implement a shared encoder architecture (plotted in blue in fig. 1 (c)), which is similar to the architecture employed in the original U-Net framework [33]. Unlike separate architecture, the shared encoder provide shared information across all timesteps, preventing models from overfitting (see section 5.3 for a discussion).
  • Separate decoders for different time intervals. Motivated by the multi-head structure introduced in the Mask Region-based Convolutional Neural Networks (Mask-RCNN) method [34], we propose to use multiple distinct decoders (plotted in colors for different intervals in fig. 1(c)), where each decoder is tailored to a specific timestep interval. The architecture of each decoder closely resembles the one utilized in [2], with deliberate adjustments made to the embedding dimensions to optimize performance.
As we observe, the primary difference in the architecture resides within the decoder structure. Intuitively, we use a decoder with fewer number of parameters for intervals closer to the noise, because the learning task is easier. We use a decoder with a larger number of parameters for intervals closer to the image.
@@ 4.2 Optimal Denoiser-based Timestep Clustering
Next, we discuss how we principally choose the interval partition time points in practice. For simplicity, we focus on the case where we partition the time $[0,1]$ into three intervals $[0,t_{1}),[t_{1},t_{2}),[t_{2},1]$, and we develop a timestep clustering method to choose the optimal $t_{1}$ and $t_{2}$. Of course, our method can be generalized to multi-stage networks with arbitrary interval numbers. However, in practice, we find that the choice of three intervals strikes a good balance between effectiveness and complexity; see our ablation study in section B.5.
To partition the time interval, we employ the optimal denoiser established in Proposition 1.
@@
1:Input: Total samples $K$, optimal denoiser function $\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x},t)$, thresholds $\alpha$, $\eta$, dataset $p_{\text{data}}$, $\mathcal{S}_{0}=\mathcal{S}_{1}=\emptyset$
@@
3:for $k\in\{1,\dots,K\}$ do
4:    $\bm{y}_{k}\sim p_{\text{data}},\bm{\epsilon}_{k}\sim\mathcal{N}(0,\mathbf{I}),t_{k}\sim[0,1]$
5:    $\mathcal{S}^{k}_{0}\leftarrow\mathcal{D}(\bm{\epsilon}^{*}_{t_{k}},\bm{\epsilon}^{*}_{0},\bm{y}_{k},\bm{\epsilon}_{k}),\mathcal{S}^{k}_{1}\leftarrow\mathcal{D}(\bm{\epsilon}^{*}_{t_{k}},\bm{\epsilon}^{*}_{1},\bm{y}_{k},\bm{\epsilon}_{k})$
6:    $\mathcal{S}_{0}\leftarrow\mathcal{S}_{0}\cup\{\left(t_{k},\mathcal{S}^{k}_{0}\right)\},\mathcal{S}_{1}\leftarrow\mathcal{S}_{1}\cup\{\left(t_{k},\mathcal{S}^{k}_{1}\right)\}$
7:end for
8:$t_{1}=\operatorname*{\text{arg~{}max}}\limits_{\tau}\left\{\tau~{}\Big{|}~{}\frac{\sum_{(t_{k},\mathcal{S}^{k}_{0})\in\mathcal{S}_{0}}\left[\mathcal{S}^{k}_{0}\cdot\mathds{1}(t_{k}\leq\tau)\right]}{\sum_{(t_{k},\mathcal{S}^{k}_{0})\in\mathcal{S}_{0}}\left[\mathds{1}(t_{k}\leq\tau)\right]}\geq\alpha\right\}$
9:$t_{2}=\operatorname*{\text{arg~{}min}}\limits_{\tau}\left\{\tau~{}\Big{|}~{}\frac{\sum_{(t_{k},\mathcal{S}^{k}_{1})\in\mathcal{S}_{0}}\left[\mathcal{S}^{k}_{1}\cdot\mathds{1}(t_{k}\geq\tau)\right]}{\sum_{(t_{k},\mathcal{S}^{k}_{1})\in\mathcal{S}_{0}}\left[\mathds{1}(t_{k}\geq\tau)\right]}\geq\alpha\right\}$
@@
where $\bm{x}_{0}\sim p_{\text{data}}(\bm{x})=\frac{1}{N}\sum_{i=1}^{N}\delta(\bm{x}-\bm{y}_{i})$, $\bm{\epsilon}\sim\mathcal{N}(0,\textbf{I})$, and $\bm{x}_{t}\sim p_{t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I})$ with perturbation parameters $s_{t}$ and $\sigma_{t}$ defined in eq. 1. Then, the optimal denoiser at $t$, defined as $\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)=\operatorname*{\text{arg~{}min}}_{\bm{\epsilon}_{\bm{\theta}}}\mathcal{L}(\bm{\epsilon}_{\bm{\theta}};t)$, is given by
$\bm{\epsilon}^{*}_{\bm{\theta}}(\bm{x};t)=\frac{1}{s_{t}\sigma_{t}}\left[\bm{x}-s_{t}\frac{\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})\bm{y}_{i}}{\sum_{i=1}^{N}\mathcal{N}(\bm{x};s_{t}\bm{y}_{i},s_{t}^{2}\sigma_{t}^{2}\textbf{I})}\right]\>.$ (4)
The proof is provided in appendix A, and the result can be generalized from recent work of Karras et al. [18], extending from a specific kernel $p_{t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};\bm{x}_{0},\sigma_{t}^{2}\textbf{I})$ to encompassing a broader scope of noise perturbation kernels, given by $p_{t}(\bm{x}_{t}|\bm{x}_{0})=\mathcal{N}(\bm{x}_{t};s_{t}\bm{x}_{0},s_{t}^{2}\sigma_{t}^{2}\textbf{I})$. For brevity, we simplify the notation of the optimal denoiser $\bm{\epsilon}^{*}_{\theta}(\bm{x},t)$ in Proposition 1 as $\bm{\epsilon}^{*}_{t}(\bm{x})$.